# Uncovering the Fragility of Trustworthy LLMs through Chinese Textual Ambiguity

### v1

@@ 1. Introduction
Large Language Models (LLMs) have demonstrated strong language understanding capabilities and are widely deployed across a range of real-world applications (Li et al., 2023; Thirunavukarasu et al., 2023; Raiaan et al., 2024). They are used to process complex instructions in multi-turn dialogues and are integrated into various systems as agents or components of AI workflows. However, LLMs still exhibit inherent limitations in trustworthiness, such as hallucinations (Azamfirei et al., 2023; Huang et al., 2025c), misunderstanding (Liu et al., 2023), and misalignment (Dung, 2023) that are particularly critical in safety-sensitive scenarios. Researchers have also invested significant effort in improving alignment, developing guardrails (Ayyamperumal and Ge, 2024), and enhancing uncertainty understanding (He et al., 2023; Shorinwa et al., [n. d.]) to enable more reliable use of LLMs.
@@ 3.2.1. Pretrained Transformer-based Text Classifier
Transformer-based pre-trained language models, such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and XLNet (Yang et al., 2019), have demonstrated strong performance across a wide range of text classification tasks. Given their effectiveness in sentence-level modeling (Tian et al., 2020; Zhang et al., 2021b, 2024c) and passage-level discrimination (Chang et al., 2020; Huang et al., 2025a) in applied classification scenarios, we adopt Transformer-based classifiers as our foundation. As a baseline, we used the pre-trained language model hfl/chinese-roberta-wwm-ext (Cui et al., 2020) as the classifier backbone. This model is based on the RoBERTa architecture and has been specifically optimized for Chinese, achieving strong performance in various NLP tasks. We further fine-tune the model with a binary classification objective to distinguish between ambiguous and unambiguous sentences.
@@ 4.2. Probing Ambiguity via Clarification Questioning
Figure 3 illustrates the step-by-step process by which an LLM addresses semantic ambiguity in an NLI scenario. When given an ambiguous premise, the model may fail to make a definitive inference judgment. It generates a clarification question to explicitly resolve the ambiguity. Once the user provides a disambiguating answer, the model determines the inference relation: entailment, contradiction, or neutral. This process provides a way to evaluate whether the model has correctly identified and understood the ambiguity. By leveraging joint reasoning to identify the minimal conditions needed for clarification or decision-making, our analysis method is also conceptually similar with existing work on explanation through factual and counterfactual analysis (Chen et al., [n. d.]).
@@ 5. Related Work
Disambiguation has been an extensively studied research topic in NLP, as ambiguity is inherently present in human language and communication. Traditional machine learning-based NLP approaches have primarily focused on word sense disambiguation (Navigli, 2009; Bevilacqua et al., 2021), employing knowledge-based methods, vector-based 1-nn classifiers, token taggers, and sequence taggers to resolve lexical ambiguity. Ambiguity detection has also been thoroughly explored in the literature. (Gleich et al., 2010) developed a taxonomy for classifying ambiguity and created POS-based and rule-based tools to detect ambiguity in requirement documents. (Ferrari and Esuli, 2019) trained word embeddings on domain-specific corpora and compared cross-domain term representations to automatically identify semantic ambiguities. Large Language Models (LLMs) demonstrate exceptional capabilities in natural language understanding and reasoning tasks. Compared to early transformer models, LLMs exhibit superior performance and flexibility in comprehending and solving multiple-choice questions across diverse subjects including history, science, and mathematics, as demonstrated on benchmarks such as MMLU (Hendrycks et al., 2020), MMLU-Pro (Wang et al., 2024), GPQA (Rein et al., 2024), and AIME (Balunović et al., 2025). LLMs also excel across multiple dimensions of language understanding, including commonsense reasoning (Krause and Stolzenburg, 2023) and interpretation of abstract concepts (Zhang et al., 2024a), and they even extend beyond the natural language domain, supporting tasks such as coding (Zhang et al., 2023), recommendation (Wu et al., 2024), forecasting, and anomaly detection (Su et al., 2024). However, existing reviews (Huang et al., 2025b; Yi et al., 2025) indicate that although LLMs demonstrate strong performance in language understanding tasks, they remain limited in their ability to capture fine-grained semantic nuances. Nevertheless, ambiguity remains a fundamental linguistic phenomenon that cannot be entirely overcome and has garnered significant attention from the research community. (Liu et al., 2023) presents an early work identifying limitations of LLMs in ambiguity understanding, and proposes AMBIENT, an English benchmark of ambiguous sentences. (Shi et al., 2025) specifically investigated ambiguity handling in questions to enhance LLM performance when confronted with ambiguous inputs. (Keluskar et al., 2024) explored improvements in LLM ambiguity handling for open-world question answering through simple prompt rewriting and context augmentation. (Mehrparvar and Pezzelle, 2024) examined ambiguity detection mechanisms in LLMs. CLAMBER (Zhang et al., 2024b) addresses ambiguity challenges in query intention understanding and information clarification requirements for LLMs in retrieval tasks. Although these studies focus on enhancing LLM performance in specific applications, they do not examine the fundamental language understanding behaviors of LLMs when processing ambiguous content. In this work, we use Chinese as a case study to investigate how LLMs encounter and handle ambiguity with specific scenes, thereby providing meaningful insights for future research on ambiguity processing in LLMs.
@@ A.1. Choice of Ambiguity Detection Model
To better handle this Chinese-specific task, we consider using language models that are pretrained on Chinese corpora and tasks (Zhang et al., 2021a; Sun et al., 2019; Cui et al., 2020). Among them, we select hfl/chinese-roberta-wwm-ext (Cui et al., 2020), a RoBERTa-based model specifically designed for Chinese. Unlike standard BERT models that apply subword-level masking, this model adopts whole-word masking (WWM), meaning that it masks entire Chinese words during pretraining. Since Chinese words often consist of multiple characters, WWM enables the model to learn more meaningful word-level representations. This property is particularly beneficial for identifying sentence-level ambiguity, where subtle differences in phrasing can lead to different interpretations.

### v2

@@ 1. Introduction
Large Language Models (LLMs) have demonstrated strong language understanding capabilities and are widely deployed across a range of real-world applications (Li et al., 2023; Thirunavukarasu et al., 2023; Raiaan et al., 2024). They are used to process complex instructions in multi-turn dialogues and are integrated into various systems as agents or components of AI workflows. However, LLMs still exhibit inherent limitations in trustworthiness, such as hallucinations (Azamfirei et al., 2023; Huang et al., 2025a), misunderstanding (Liu et al., 2023), and misalignment (Dung, 2023) that are particularly critical in safety-sensitive scenarios. Researchers have also invested significant effort in improving alignment, developing guardrails (Ayyamperumal and Ge, 2024), and enhancing uncertainty understanding (He et al., 2023; Shorinwa et al., ) to enable more reliable use of LLMs.
@@ 3.2.1. Pretrained Transformer-based Text Classifier
Transformer-based pre-trained language models, such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and XLNet (Yang et al., 2019), have demonstrated strong performance across a wide range of text classification tasks. Given their effectiveness in sentence-level modeling (Tian et al., 2020; Zhang et al., 2021a, 2024c) and passage-level discrimination (Chang et al., 2020; Huang et al., 2025b) in applied classification scenarios, we adopt Transformer-based classifiers as our foundation. As a baseline, we used the pre-trained language model hfl/chinese-roberta-wwm-ext (Cui et al., 2020) as the classifier backbone. This model is based on the RoBERTa architecture and has been specifically optimized for Chinese, achieving strong performance in various NLP tasks. We further fine-tune the model with a binary classification objective to distinguish between ambiguous and unambiguous sentences.
@@ 4.2. Probing Ambiguity via Clarification Questioning
Figure 3 illustrates the step-by-step process by which an LLM addresses semantic ambiguity in an NLI scenario. When given an ambiguous premise, the model may fail to make a definitive inference judgment. It generates a clarification question to explicitly resolve the ambiguity. Once the user provides a disambiguating answer, the model determines the inference relation: entailment, contradiction, or neutral. This process provides a way to evaluate whether the model has correctly identified and understood the ambiguity. By leveraging joint reasoning to identify the minimal conditions needed for clarification or decision-making, our analysis method is also conceptually similar with existing work on explanation through factual and counterfactual analysis (Chen et al., ).
@@ 5. Related Work
Disambiguation has been an extensively studied research topic in NLP, as ambiguity is inherently present in human language and communication. Traditional machine learning-based NLP approaches have primarily focused on word sense disambiguation (Navigli, 2009; Bevilacqua et al., 2021), employing knowledge-based methods, vector-based 1-nn classifiers, token taggers, and sequence taggers to resolve lexical ambiguity. Ambiguity detection has also been thoroughly explored in the literature. (Gleich et al., 2010) developed a taxonomy for classifying ambiguity and created POS-based and rule-based tools to detect ambiguity in requirement documents. (Ferrari and Esuli, 2019) trained word embeddings on domain-specific corpora and compared cross-domain term representations to automatically identify semantic ambiguities. Large Language Models (LLMs) demonstrate exceptional capabilities in natural language understanding and reasoning tasks. Compared to early transformer models, LLMs exhibit superior performance and flexibility in comprehending and solving multiple-choice questions across diverse subjects including history, science, and mathematics, as demonstrated on benchmarks such as MMLU (Hendrycks et al., 2020), MMLU-Pro (Wang et al., 2024), GPQA (Rein et al., 2024), and AIME (Balunović et al., 2025). LLMs also excel across multiple dimensions of language understanding, including commonsense reasoning (Krause and Stolzenburg, 2023) and interpretation of abstract concepts (Zhang et al., 2024b), and they even extend beyond the natural language domain, supporting tasks such as coding (Zhang et al., 2023), recommendation (Wu et al., 2024), forecasting, and anomaly detection (Su et al., 2024). However, existing reviews (Huang et al., 2025c; Yi et al., 2025) indicate that although LLMs demonstrate strong performance in language understanding tasks, they remain limited in their ability to capture fine-grained semantic nuances. Nevertheless, ambiguity remains a fundamental linguistic phenomenon that cannot be entirely overcome and has garnered significant attention from the research community. (Liu et al., 2023) presents an early work identifying limitations of LLMs in ambiguity understanding, and proposes AMBIENT, an English benchmark of ambiguous sentences. (Shi et al., 2025) specifically investigated ambiguity handling in questions to enhance LLM performance when confronted with ambiguous inputs. (Keluskar et al., 2024) explored improvements in LLM ambiguity handling for open-world question answering through simple prompt rewriting and context augmentation. (Mehrparvar and Pezzelle, 2024) examined ambiguity detection mechanisms in LLMs. CLAMBER (Zhang et al., 2024a) addresses ambiguity challenges in query intention understanding and information clarification requirements for LLMs in retrieval tasks. Although these studies focus on enhancing LLM performance in specific applications, they do not examine the fundamental language understanding behaviors of LLMs when processing ambiguous content. In this work, we use Chinese as a case study to investigate how LLMs encounter and handle ambiguity with specific scenes, thereby providing meaningful insights for future research on ambiguity processing in LLMs.
@@ A.1. Choice of Ambiguity Detection Model
To better handle this Chinese-specific task, we consider using language models that are pretrained on Chinese corpora and tasks (Zhang et al., 2021b; Sun et al., 2019; Cui et al., 2020). Among them, we select hfl/chinese-roberta-wwm-ext (Cui et al., 2020), a RoBERTa-based model specifically designed for Chinese. Unlike standard BERT models that apply subword-level masking, this model adopts whole-word masking (WWM), meaning that it masks entire Chinese words during pretraining. Since Chinese words often consist of multiple characters, WWM enables the model to learn more meaningful word-level representations. This property is particularly beneficial for identifying sentence-level ambiguity, where subtle differences in phrasing can lead to different interpretations.